OpenStudy (j.f):

Medal x is a prime number less than 12. What is the probability that 2≤ x≤ 6? I'm not understanding how you would find probability

OpenStudy (anonymous):

How many prime numbers less than 12 lie between 2 and 6 (inclusive)?

OpenStudy (j.f):

3 and 5

OpenStudy (anonymous):

You're missing one... Notice that \(2\le x\) means 2 is also considered. Now how many prime numbers are there that are less than 12?

OpenStudy (j.f):

5 numbers less than 12

OpenStudy (anonymous):

Right. So the probability must be \(\dfrac{3}{5}\), correct? We're fixing \(x\) to be a prime less than 12, and only 3 of those numbers satisfy the condition that \(2\le x\le6\).

OpenStudy (j.f):

That makes total sense